### v1

# New Linear Model of a Composite Energy Storage System with Realizable Dispatch Guarantees

Abstract
To optimize battery dispatch, a model is required that can predict the state of charge (SOC) trajectory and ensure dispatch is admissible (i.e., does not lead to unexpected SOC saturation). But battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint. In this paper, we consider a composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery. This permits convex optimization of the composite battery SOC trajectory while ensuring admissibility of the resulting (aggregated) power schedule and disaggregation to the individual energy storage elements.
Index Terms:
Battery, energy storage, optimization, convex.
I Introduction
Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s transition to renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation]. Grid-scale BESSs are comprised of several battery racks that can be split into several battery modules and further into battery cells. When optimizing BESS dispatch the individual elements are lumped as one composite battery and the composite dispatch is shared equally between the battery elements. Optimizing composite dispatch is a non-convex problem since a complementarity constraint is needed to avoid simultaneous charging and discharging. The complementarity constraint can be formulated with a non-linear constraint, e.g., $P_{\text{c}}[k]P_{\text{d}}[k]=0$, or by the use of binary variables. To obtain a convex formulation, the complementarity constraint is typically omitted. The optimal power trajectories then exhibit simultaneous charging and discharging, but only the net power trajectory is applied. Implementing the net power trajectory results in a discrepancy between the predicted optimal SOC trajectory and the realized SOC trajectory [RobustNawaf, D.Pozo_paper]. This discrepancy can lead to unexpected SOC saturation, resulting in un-realizable power trajectories. In [RobustNawaf, Mazen_CDC_paper], a robust linear battery model is presented as an inner approximation that guarantees that any optimal input power trajectory is realizable, but it is conservative over longer prediction horizons. Previous work has always assumed equal power-sharing between battery elements is enforced. This work revisits battery modeling by considering a composite battery made of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by disaggregation to the individual elements without the equal power sharing paradigm. The paper characterizes the set of realizable composite power trajectories that can be disaggregated to the elements while ensuring element-wise power limits, energy limits, and complementarity constraints are satisfied. A priority-based strategy is employed to perform disaggregation.
II Composite Battery Model
A general BESS is shown in Fig 1, where $P_{\text{c/d}}[k]$, $\eta_{\text{c/d}}$, and $E[k]$ are the composite charging/discharging power, charge/discharge efficiency and the SOC at timestep $k$, respectively.
Figure 1: Example of a composite battery with dc-coupled battery elements.
The BESS shown in Fig 1 is a composite of $N$ identical elements. Examples of this are DC-coupled battery racks connected to form a single BESS unit, or even multiple AC-coupled BESS units coordinated to form an aggregated (virtual) battery. For element $i\in\{1,\ldots,N\}$, the charge/discharge power and the SOC are denoted by $P^{i}_{\text{c/d}}[k]$ and $E^{i}[k]$, respectively.
A BESS dispatch needs to ensure that power limits, energy limits, and complementarity constraints are satisfied for all elements
Battery element $i$
$\begin{split}&E^{i}[k+1]=E^{i}[k]+\\ &\hskip 2.84526pt\Delta t\eta_{\text{c}}P_{\text{c}}^{i}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}[k]\end{split}$ (1a)
$E^{i}[0]=E^{i}_{0}$ (1b)
$0=P_{\text{c}}^{i}[k]P_{\text{d}}^{i}[k]$ (1c)
$0\leq E^{i}[k]\leq{E}_{\text{max}}$ (1d)
$0\leq P_{\text{c/d}}^{i}[k]\leq{P}_{\text{c/d,max}}$ (1e)
Balanced composite battery
$\begin{split}&E[k+1]=E[k]+\\ &\hskip 2.84526pt\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]\end{split}$ (2a)
$E[0]=NE_{0}$ (2b)
$0=P_{\text{c}}[k]P_{\text{d}}[k]$ (2c)
$0\leq E[k]\leq N{E}_{\text{max}}$ (2d)
$0\leq P_{\text{c/d}}[k]\leq N{P}_{\text{c/d,max}}$ (2e)
$\forall i=1,\ldots,N$ and $\forall k=0,\ldots,K-1$, where $\Delta t$ is the time step. Note, it is assumed that the individual elements are identical with the same power and energy limits, ${P}_{\text{c/d,max}}$ and ${E}_{\text{max}}$. This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$, as shown in (2). Typically, the composite battery system maintains balanced SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and equal power sharing dispatch across its elements, which means that a feasible composite dispatch implicitly satisfies all elements’ constraints in (1). Under equal power sharing, a single complementarity constraint at the composite battery level ($P_{\text{c}}[k]P_{\text{d}}[k]=0$) is sufficient to satisfy (1c) for all $i$.
However, under unequal power sharing, this does not hold. In fact, with unequal power sharing, we can omit the complementarity constraint at the composite level, but then have to disaggregate the composite power trajectory across the $N$ elements while satisfying (1c). That is, unequal power sharing increases capability of the composite battery, but at the cost of computational complexity with $N$ complementary constraints. Thus, this paper presents a priority-based disaggregation strategy that engenders a provably convex set of realizable composite power trajectories.
III Composite Trajectory Disaggregation
A control policy is used to disaggregate a desired piecewise constant composite power inputs $P_{\text{c/d}}[k]$ into $N$ individual element charge/discharge inputs, $\{P^{i}_{\text{c/d}}[l]\}_{i=1}^{N}$. The controller operates on a faster timescale with $L=MK$ timesteps and $M\geq 1$ and ensures that for each time-step on the fast timescale, $l\in\{0,\ldots,L-1\}$, we have
$\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k],$ (3)
where $k=\lfloor\frac{l}{M}\rfloor$. One such controller is the presented priority stack controller discussed below. However, the priority stack controller does not guarantee that constraints in (1) are met, even if the composite power trajectory satisfies the composite power and energy limits. Specifically, under the priority stack policy, (1c) and (1d) may be violated. However, this paper develops sufficient conditions that with the priority stack controller enables characterization of $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ that guarantees (1) is satisfied.
III-A Priority Stack Controller (PSC)
Based on the SOC, the priority stack controller (PSC) sorts all $N$ elements every (fast) time-step $l$. The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority. When selecting the highest-priority element for charging/discharging that element’s power is maximized, i.e., $P^{i}_{\text{c/d}}(l)=P_{\text{c/d,max}}$, before actuating another charging/discharging element from the stack. This process is repeated every time-step $l$, adapting to changing element SOCs. The priority stack results in $N_{\text{c}}$ elements with lowest SOCs charging and $N_{\text{d}}$ elements with highest SOCs discharging, as follows:
$N_{\text{c}}(l):=\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\quad\land\quad N_{\text{d}}(l):=\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}.$ (4)
The PSC is summarized in Algorithm 1 and begets Property 1.
Algorithm 1 Priority Stack Controller (PSC)
1:Given $P^{\ast}_{\text{c}}[k],P^{\ast}_{\text{d}}[k]$ subject to (2)
2:for each timestep $l$ do
3:    Sort elements based on $E^{i}(l)$ in ascending order
4:    Calculate $N_{\text{c/d}}(l):=\lceil\frac{P^{\ast}_{\text{c/d}}[k]}{P_{\text{c/d,max}}}\rceil$
5:    Assign $P^{i}_{\text{c/d}}(l)$ for first $N_{\text{c}}(l)$/last $N_{\text{d}}(l)$ elements
6:    Update SOC for all elements: $E^{i}(l)\rightarrow E^{i}(l+1)$
7:end for
Property 1.
Under PSC, if elements $i,j$ satisfy $E^{i}(l)\geq E^{j}(l)$ , then $P^{i}_{\text{c}}(l)\leq P^{j}_{\text{c}}(l)$ and $P^{i}_{\text{d}}(l)\geq P^{j}_{\text{d}}(l)$.
Next, we analyze the effects of the PSC on (1), on complementarity conditions (1c), and on energy limits (1d).
III-B Guaranteeing Element Complementarity Constraints
To ensure the complementarity constraints are satisfied at the element level, $(P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)=0\quad\forall i,\forall l)$, the priority stack cannot issue a charging and discharging demand to the same battery element $i$ simultaneously. Thus, any overlap in the selection of $N_{\text{c}}(l)$ charging and $N_{\text{d}}(l)$ discharging elements imply a violation of the complementarity constraints, which must, therefore, satisfy:
$N_{\text{c}}(l)+N_{\text{d}}(l)\leq N\quad\forall l.$ (5)
However, considering (4), we can tighten (5) as follows:
$\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\leq N-1.$ (6)
The feasible region dictated by (5) and (6) is shown in Fig. 2.
Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right).
Proposition 1.
If $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy (6) then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (5).
Proof.
Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c/d}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$.
Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (6), we get
$\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}\leq\Big{\lceil}N-1-(\Gamma_{\text{c}}+\Delta_{\text{c}})\Big{\rceil}=N-1-\Gamma_{\text{c}}.$ (7)
Adding inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$, satisfying (5). ∎
Thus, adding (6) to the composite formulation in (2) will satisfy element complementarity constraints under PSC. Similarly, we want to relate element energy limits to those in the composite formulation.
III-C Relating Element and Composite Energy Limit Constraints
Under PSC, the battery elements are charged/discharged individually every $l$, which results in different SOC trajectories. The spread of element SOCs can impact realizability of the composite dispatch especially for larger element SOC spread. This complexity can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits, which tightens the composite energy limits as
$N\Delta E_{\text{max}}\leq E[k]\leq N(E_{\text{max}}-\Delta E_{\text{max}}),$ (8)
where $\Delta E_{\text{max}}$ is given by
$\Delta E_{\text{max}}:=\frac{\Delta t}{M}\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right).$ (9)
Proposition 2.
A composite battery dispatch that satisfies (8), will not violate any individual element’s energy limit under PSC, provided that $\Delta E(0)\leq\Delta E_{\text{max}}$.
Proof.
Given element $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
$E^{i}(l+1)-E^{j}(l+1)$ $=\left(E^{i}(l)-E^{j}(l)\right)+$ (10)
$\frac{\Delta t}{M}\eta_{\text{c}}\left(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\right)-\frac{\Delta t}{M\eta_{\text{d}}}\left(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\right).$
From Property 1 and power limits, we have $-P_{\text{c,max}}\leq(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l))\leq 0$ and $0\leq(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l))\leq P_{\text{d,max}}$, which implies
$E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l)$ (11a)
and
$E^{i}(l+1)-E^{j}(l+1)\geq$ $-\frac{\Delta t}{M}\left(\eta_{\text{c}}P_{\text{c,max}}+\frac{P_{\text{d,max}}}{\eta_{\text{d}}}\right)=-\Delta E_{\text{max}},$ (11b)
where the last inequality in (11b) uses $E^{i}(l)-E^{j}(l)\geq 0$. Now, we can bound $\Delta E(l+1)\leq\max\left(\Delta E(l),\Delta E_{\text{max}}\right)=\Delta E_{\text{max}}.$ Therefore, $\Delta E(l)\leq\Delta E_{\text{max}}\Rightarrow\Delta E(l+1)\leq\Delta E_{\text{max}}$, which implies that for a composite battery with initial element SOC $E^{i}(0)$, if $\Delta E(0)\leq\Delta E_{\text{max}}$, then $\Delta E(l)\leq\Delta E_{\text{max}}\forall l=0,\ldots,MK$
III-D Realizable Composite Power Trajectories
The following theorem provides a linear formulation for realizable dispatch of a composite battery.
Theorem 1.
If $\Delta E[0]\leq\Delta E_{\text{max}}$, then any feasible composite battery dispatch schedule $\{P_{\text{c/d}}^{\ast}[k]\}_{k=0}^{K-1}$ resulting from optimizing over (12) will satisfy (1) under PSC and is realizable across $N$ elements.
[…]
$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}$ $+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq\frac{N-1}{N}$ (12b)
$N\Delta E_{\text{max}}$ $\leq E[k]\leq N\left(E_{\text{max}}-\Delta E_{\text{max}}\right)$ (12c)
Proof.
From Proposition 1, (12b) ensures satisfaction of (1c). Similarly, with $\Delta E[0]\leq\Delta E_{\text{max}}$, Proposition 2 guarantees that (1d) holds for all $N$ elements. Thus, any feasible composite dispatch is realizable by its $N$ elements.∎
Note that for a fixed composite battery, if the elements’ $P_{\text{c/d,max}}$ decreases (i.e., $N$ increases), the cutting plane in (12b) approaches the cutting plane of the relaxed model. Furthermore, for $M\gg 1$, $\Delta E_{\text{max}}$ tends to zero. That is, with faster control and more granular elements, (12) approaches the relaxed battery model defined by (12a); however, with guaranteed realizability of any feasible dispatch.
IV Simulation Results
We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}[k]}\right)^{2}$. For comparison, relaxed and robust models from [RobustNawaf] are used with equal power-sharing enforced. Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison. The chosen element battery parameters are obtained from a Tesla Powerwall [tesla_2019] and we set $N=1000$ and $M=1$. The results from using each model are shown in Fig. 3.
Figure 3: Reference tracking optimization.
The RCB model and relaxed battery model produce the same optimal dispatch. However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate and cause the sudden drop in power seen in Fig. 3. The MILP, robust, and RCB models produced realizable results, with the robust formulation resulting in worse tracking performance. The MILP and RCB formulations produced comparable results, however, with a significant computational advantage for the RCB model.
V Conclusion
A new convex dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging. A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied. Previous battery optimization implicitly assumed equal power sharing between all elements. By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models, which yield realizable power trajectories.
[…]

### v2

@@
-New Linear Model of a Composite Energy Storage System with Realizable Dispatch Guarantees
+Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees
 Abstract
 To optimize battery dispatch, a model is required that can predict the state of charge (SOC) trajectory and ensure dispatch is admissible (i.e., does not lead to unexpected SOC saturation).
-But battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint.
+However, battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint.
 In this paper, we consider a composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery.
@@
 I Introduction
-Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s transition to renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation].
+Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s transition to renewable energy and improve reliability  [Battery_review_paper].
 Grid-scale BESSs are comprised of several battery racks that can be split into several battery modules and further into battery cells.
@@
 Optimizing composite dispatch is a non-convex problem since a complementarity constraint is needed to avoid simultaneous charging and discharging.
-The complementarity constraint can be formulated with a non-linear constraint, e.g., $P_{\text{c}}[k]P_{\text{d}}[k]=0$, or by the use of binary variables.
+The complementarity constraint can be formulated with a non-linear constraint, or by the use of binary variables.
 To obtain a convex formulation, the complementarity constraint is typically omitted.
@@
 Previous work has always assumed equal power-sharing between battery elements is enforced.
-This work revisits battery modeling by considering a composite battery made of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by disaggregation to the individual elements without the equal power sharing paradigm.
+This work revisits battery modeling by considering a composite battery made of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements.
 The paper characterizes the set of realizable composite power trajectories that can be disaggregated to the elements while ensuring element-wise power limits, energy limits, and complementarity constraints are satisfied.
@@
 II Composite Battery Model
-A general BESS is shown in Fig 1, where $P_{\text{c/d}}[k]$, $\eta_{\text{c/d}}$, and $E[k]$ are the composite charging/discharging power, charge/discharge efficiency and the SOC at timestep $k$, respectively.
+We consider a generic BESS as shown in Fig 1.
+It is composed of $N$ identical elements.
+Examples of this are DC-coupled battery racks connected to form a single BESS unit, or even multiple AC-coupled BESS units coordinated to form an aggregated (virtual) battery.
 Figure 1: Example of a composite battery with dc-coupled battery elements.
-The BESS shown in Fig 1 is a composite of $N$ identical elements.
-Examples of this are DC-coupled battery racks connected to form a single BESS unit, or even multiple AC-coupled BESS units coordinated to form an aggregated (virtual) battery.
-For element $i\in\{1,\ldots,N\}$, the charge/discharge power and the SOC are denoted by $P^{i}_{\text{c/d}}[k]$ and $E^{i}[k]$, respectively.
-A BESS dispatch needs to ensure that power limits, energy limits, and complementarity constraints are satisfied for all elements
+For element $i$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$.
+The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOC at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively.
+A realizable BESS dispatch must satisfy power limits (${P}_{\text{c/d,max}}$), energy limits (${E}_{\text{max}}$), and complementarity constraints for all battery elements ($\forall i=1,\ldots,N$) and time steps ($\forall l=0,\ldots,L-1$).
+Note that all elements have identical power and energy limits.
+This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$.
 Battery element $i$
-$\begin{split}&E^{i}[k+1]=E^{i}[k]+\\ &\hskip 2.84526pt\Delta t\eta_{\text{c}}P_{\text{c}}^{i}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}[k]\end{split}$ (1a)
-$E^{i}[0]=E^{i}_{0}$ (1b)
-$0=P_{\text{c}}^{i}[k]P_{\text{d}}^{i}[k]$ (1c)
-$0\leq E^{i}[k]\leq{E}_{\text{max}}$ (1d)
-$0\leq P_{\text{c/d}}^{i}[k]\leq{P}_{\text{c/d,max}}$ (1e)
+$\begin{split}&E^{i}(l+1)=E^{i}(l)+\\ &\delta t\eta_{\text{c}}P_{\text{c}}^{i}(l)-\delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}(l)\end{split}$ (1a)
+$E^{i}(0)=E^{i}_{0}$ (1b)
+$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)$ (1c)
+$0\leq E^{i}(l+1)\leq{E}_{\text{max}}$ (1d)
+$0\leq P_{\text{c/d}}^{i}(l)\leq{P}_{\text{c/d,max}}$ (1e)
 Balanced composite battery
-$\begin{split}&E[k+1]=E[k]+\\ &\hskip 2.84526pt\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]\end{split}$ (2a)
-$E[0]=NE_{0}$ (2b)
+$\begin{split}&E[k+1]=E[k]+\\ &\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]\end{split}$ (2a)
+$E[0]=\sum_{i}E^{i}_{0}$ (2b)
 $0=P_{\text{c}}[k]P_{\text{d}}[k]$ (2c)
-$0\leq E[k]\leq N{E}_{\text{max}}$ (2d)
+$0\leq E[k+1]\leq N{E}_{\text{max}}$ (2d)
 $0\leq P_{\text{c/d}}[k]\leq N{P}_{\text{c/d,max}}$ (2e)
-$\forall i=1,\ldots,N$ and $\forall k=0,\ldots,K-1$, where $\Delta t$ is the time step.
-Note, it is assumed that the individual elements are identical with the same power and energy limits, ${P}_{\text{c/d,max}}$ and ${E}_{\text{max}}$.
-This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$, as shown in (2).
-Typically, the composite battery system maintains balanced SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and equal power sharing dispatch across its elements, which means that a feasible composite dispatch implicitly satisfies all elements’ constraints in (1).
-Under equal power sharing, a single complementarity constraint at the composite battery level ($P_{\text{c}}[k]P_{\text{d}}[k]=0$) is sufficient to satisfy (1c) for all $i$.
+The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$.
+We denote these scheduling time steps with index $[k]$ in square brackets, where $k=0,\ldots,K-1$ and $L=MK$.
+Constant power in step $k$ implies the following constraint for all $l,k$ with $k=\lfloor\frac{l}{M}\rfloor$:
+$\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k].$ (3)
+Typically, a composite BESS is modeled and operated in a balanced way: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and equal power sharing across its elements.
+This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (2).
+Under equal power sharing, a single complementarity constraint at the composite battery level, (2c), is sufficient to satisfy (1c) for all $i$ and all $l$.
+Under balanced operation, a feasible composite dispatch implicitly satisfies all elements’ constraints in (1).
 However, under unequal power sharing, this does not hold.
 In fact, with unequal power sharing, we can omit the complementarity constraint at the composite level, but then have to disaggregate the composite power trajectory across the $N$ elements while satisfying (1c).
-That is, unequal power sharing increases capability of the composite battery, but at the cost of computational complexity with $N$ complementary constraints.
-Thus, this paper presents a priority-based disaggregation strategy that engenders a provably convex set of realizable composite power trajectories.
+That is, unequal power sharing increases the capability of the composite battery, but at the cost of computational complexity with $N$ complementary constraints.
+Confronting this challenge, this paper presents a priority-based disaggregation strategy that engenders a provably convex linear set of realizable composite power trajectories.
+These trajectories permit simultaneous charging and discharging of the BESS and approximate the full solution space arbitrarily closely in well-defined limits.
 III Composite Trajectory Disaggregation
-A control policy is used to disaggregate a desired piecewise constant composite power inputs $P_{\text{c/d}}[k]$ into $N$ individual element charge/discharge inputs, $\{P^{i}_{\text{c/d}}[l]\}_{i=1}^{N}$.
-The controller operates on a faster timescale with $L=MK$ timesteps and $M\geq 1$ and ensures that for each time-step on the fast timescale, $l\in\{0,\ldots,L-1\}$, we have
-$\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k],$ (3)
-where $k=\lfloor\frac{l}{M}\rfloor$.
+A control policy is used to disaggregate a desired piecewise constant composite power inputs $P_{\text{c/d}}[k]$ into $N$ individual element charge/discharge inputs, $\{P^{i}_{\text{c/d}}(l)\}_{i=1}^{N}$.
 One such controller is the presented priority stack controller discussed below.
-However, the priority stack controller does not guarantee that constraints in (1) are met, even if the composite power trajectory satisfies the composite power and energy limits.
-Specifically, under the priority stack policy, (1c) and (1d) may be violated.
-However, this paper develops sufficient conditions that with the priority stack controller enables characterization of $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ that guarantees (1) is satisfied.
+Sufficient conditions are developed that – in combination with the controller – enable the characterization of $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ that guarantees (1) is satisfied.
 III-A Priority Stack Controller (PSC)
-Based on the SOC, the priority stack controller (PSC) sorts all $N$ elements every (fast) time-step $l$.
+Every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC.
 The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority.
@@
 Algorithm 1 Priority Stack Controller (PSC)
-1:Given $P^{\ast}_{\text{c}}[k],P^{\ast}_{\text{d}}[k]$ subject to (2)
+1:Given $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$
 2:for each timestep $l$ do
 3:    Sort elements based on $E^{i}(l)$ in ascending order
-4:    Calculate $N_{\text{c/d}}(l):=\lceil\frac{P^{\ast}_{\text{c/d}}[k]}{P_{\text{c/d,max}}}\rceil$
+4:    Calculate $N_{\text{c/d}}(l):=\lceil\frac{P_{\text{c/d}}[k]}{P_{\text{c/d,max}}}\rceil$
 5:    Assign $P^{i}_{\text{c/d}}(l)$ for first $N_{\text{c}}(l)$/last $N_{\text{d}}(l)$ elements
@@
 Under PSC, if elements $i,j$ satisfy $E^{i}(l)\geq E^{j}(l)$ , then $P^{i}_{\text{c}}(l)\leq P^{j}_{\text{c}}(l)$ and $P^{i}_{\text{d}}(l)\geq P^{j}_{\text{d}}(l)$.
+III-B Guaranteeing Element Complementarity Constraints
 Next, we analyze the effects of the PSC on (1), on complementarity conditions (1c), and on energy limits (1d).
-III-B Guaranteeing Element Complementarity Constraints
 To ensure the complementarity constraints are satisfied at the element level, $(P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)=0\quad\forall i,\forall l)$, the priority stack cannot issue a charging and discharging demand to the same battery element $i$ simultaneously.
@@
 $N_{\text{c}}(l)+N_{\text{d}}(l)\leq N\quad\forall l.$ (5)
-However, considering (4), we can tighten (5) as follows:
+Proposition 1.
+Under the PSC, if $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy
 $\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\leq N-1.$ (6)
-The feasible region dictated by (5) and (6) is shown in Fig. 2.
-Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right).
-Proposition 1.
-If $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy (6) then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (5).
+then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (5).
 Proof.
-Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c/d}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$.
+Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$.
 Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (6), we get
 $\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}\leq\Big{\lceil}N-1-(\Gamma_{\text{c}}+\Delta_{\text{c}})\Big{\rceil}=N-1-\Gamma_{\text{c}}.$ (7)
-Adding inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$, satisfying (5). ∎
-Thus, adding (6) to the composite formulation in (2) will satisfy element complementarity constraints under PSC.
+Adding inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$ for any $l$. ∎
+Thus, adding (6) to the composite formulation in (2) will satisfy element complementarity constraints under PSC. The feasible regions dictated by (5) and (6) are shown in Fig. 2.
+Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are not considered here.
+III-C Relating Element and Composite Energy Limit Constraints
 Similarly, we want to relate element energy limits to those in the composite formulation.
-III-C Relating Element and Composite Energy Limit Constraints
-Under PSC, the battery elements are charged/discharged individually every $l$, which results in different SOC trajectories.
-The spread of element SOCs can impact realizability of the composite dispatch especially for larger element SOC spread.
-This complexity can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits, which tightens the composite energy limits as
-$N\Delta E_{\text{max}}\leq E[k]\leq N(E_{\text{max}}-\Delta E_{\text{max}}),$ (8)
-where $\Delta E_{\text{max}}$ is given by
-$\Delta E_{\text{max}}:=\frac{\Delta t}{M}\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right).$ (9)
+Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOC trajectories.
+The spread of element SOCs can impact realizability of the composite dispatch when individual elements hit their SOC limits.
+Constraint violations can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
 Proposition 2.
-A composite battery dispatch that satisfies (8), will not violate any individual element’s energy limit under PSC, provided that $\Delta E(0)\leq\Delta E_{\text{max}}$.
+Under the PSC, a composite dispatch that satisfies
+$N\varepsilon\leq E[k]\leq N(E_{\text{max}}-\varepsilon),$ (8)
+with
+$\varepsilon:=\frac{\Delta t}{M}\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right),$ (9)
+will not violate any individual element’s energy limit, provided that $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$.
 Proof.
-Given element $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$. Using (1a) the following is obtained,
+Define the maximum difference in SOC across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOC spread).
+Consider element SOCs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$.
+Using (1a) the following is obtained,
 $E^{i}(l+1)-E^{j}(l+1)$ $=\left(E^{i}(l)-E^{j}(l)\right)+$ (10)
 $\frac{\Delta t}{M}\eta_{\text{c}}\left(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\right)-\frac{\Delta t}{M\eta_{\text{d}}}\left(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\right).$
-From Property 1 and power limits, we have $-P_{\text{c,max}}\leq(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l))\leq 0$ and $0\leq(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l))\leq P_{\text{d,max}}$, which implies
-$E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l)$ (11a)
-and
-$E^{i}(l+1)-E^{j}(l+1)\geq$ $-\frac{\Delta t}{M}\left(\eta_{\text{c}}P_{\text{c,max}}+\frac{P_{\text{d,max}}}{\eta_{\text{d}}}\right)=-\Delta E_{\text{max}},$ (11b)
-where the last inequality in (11b) uses $E^{i}(l)-E^{j}(l)\geq 0$.
-Now, we can bound $\Delta E(l+1)\leq\max\left(\Delta E(l),\Delta E_{\text{max}}\right)=\Delta E_{\text{max}}.$ Therefore, $\Delta E(l)\leq\Delta E_{\text{max}}\Rightarrow\Delta E(l+1)\leq\Delta E_{\text{max}}$, which implies that for a composite battery with initial element SOC $E^{i}(0)$, if $\Delta E(0)\leq\Delta E_{\text{max}}$, then $\Delta E(l)\leq\Delta E_{\text{max}}\forall l=0,\ldots,MK$
+From Property 1 and (1e), we have $-P_{\text{c,max}}\leq P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\leq 0$ and $0\leq P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\leq P_{\text{d,max}}$, which implies
+$E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l),$ (11a)
+$E^{i}(l+1)-E^{j}(l+1)\geq$ $-\frac{\Delta t}{M}\left(\eta_{\text{c}}P_{\text{c,max}}+\frac{P_{\text{d,max}}}{\eta_{\text{d}}}\right)=-\varepsilon,$ (11b)
+where the inequality in (11b) uses $E^{i}(l)-E^{j}(l)\geq 0$.
+Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$.
+Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$, we maintain $\Delta E(l)\leq\varepsilon$ for all $l$.
+Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (8) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$.
+Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOC.
+Because the SOC spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$.
+Combined, these inequalities show that (1d) holds for all $i$ and all $l$.
+∎
 III-D Realizable Composite Power Trajectories
@@
 Theorem 1.
-If $\Delta E[0]\leq\Delta E_{\text{max}}$, then any feasible composite battery dispatch schedule $\{P_{\text{c/d}}^{\ast}[k]\}_{k=0}^{K-1}$ resulting from optimizing over (12) will satisfy (1) under PSC and is realizable across $N$ elements.
+If $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, then any feasible composite battery dispatch schedule $\{P_{\text{c}}[k],P_{\text{d}}[k]\}_{k=0}^{K-1}$ satisfying
 […]
-$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}$ $+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq\frac{N-1}{N}$ (12b)
-$N\Delta E_{\text{max}}$ $\leq E[k]\leq N\left(E_{\text{max}}-\Delta E_{\text{max}}\right)$ (12c)
+$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}$ $\leq\frac{N-1}{N}$ $\forall k=0,\ldots,K-1$ (12b)
+$N\varepsilon\leq E[k]$ $\leq N\left(E_{\text{max}}-\varepsilon\right)$ $\forall k=0,\ldots,K$ (12c)
+$P_{\text{c}}[k],P_{\text{d}}[k]$ $\geq 0$ $\forall k=0,\ldots,K-1,$ (12d)
+will satisfy (1) for all $N$ elements under PSC and is, thus, realizable.
 Proof.
+Eqs.
+(1a) and (3) imply (2a) and (1b) implies (2b).
 From Proposition 1, (12b) ensures satisfaction of (1c).
-Similarly, with $\Delta E[0]\leq\Delta E_{\text{max}}$, Proposition 2 guarantees that (1d) holds for all $N$ elements.
-Thus, any feasible composite dispatch is realizable by its $N$ elements.∎
-Note that for a fixed composite battery, if the elements’ $P_{\text{c/d,max}}$ decreases (i.e., $N$ increases), the cutting plane in (12b) approaches the cutting plane of the relaxed model.
-Furthermore, for $M\gg 1$, $\Delta E_{\text{max}}$ tends to zero.
-That is, with faster control and more granular elements, (12) approaches the relaxed battery model defined by (12a); however, with guaranteed realizability of any feasible dispatch.
+Similarly, with $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, Proposition 2 guarantees that (1d) holds, and the PSC with (12d) and (3) ensure that (1e) holds for all $N$ elements.
+Thus, any feasible composite dispatch is realizable.∎
+Note that for a fixed composite battery, if $N$ increases, the cutting plane in (12b) approaches the cutting plane of the relaxed model.
+Furthermore, for $M\gg 1$, $\varepsilon$ in (9) tends to zero.
+That is, with faster control and more granular elements, (12) approaches the relaxed battery model defined in [Mazen_CDC_paper]; however, with guaranteed realizability.
 IV Simulation Results
-We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}[k]}\right)^{2}$.
-For comparison, relaxed and robust models from [RobustNawaf] are used with equal power-sharing enforced.
+We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}$.
+For comparison, relaxed and robust models from [RobustNawaf] are used under balanced operation.
 Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison.
-The chosen element battery parameters are obtained from a Tesla Powerwall [tesla_2019] and we set $N=1000$ and $M=1$.
+The chosen element battery parameters are obtained from a Tesla Powerwall and we set $N=1000$ and $M=1$.
 The results from using each model are shown in Fig.
@@
 Figure 3: Reference tracking optimization.
-The RCB model and relaxed battery model produce the same optimal dispatch.
-However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate and cause the sudden drop in power seen in Fig.
+The RCB model and relaxed battery model produced similar optimal dispatch.
+However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate, resulting in the sudden drop in power seen in Fig.
 3.
@@
 V Conclusion
-A new convex dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging.
+A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging.
 A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied.
 Previous battery optimization implicitly assumed equal power sharing between all elements.
-By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models, which yield realizable power trajectories.
+By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories.
+Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems.
 […]

### v4

@@
 Abstract
-To optimize battery dispatch, a model is required that can predict the state of charge (SOC) trajectory and ensure dispatch is admissible (i.e., does not lead to unexpected SOC saturation).
+To optimize the dispatch of batteries, a model is required that can predict the state of charge (SOC) trajectory for a chosen open-loop power schedule to ensure admissibility (i.e., that schedule can be realized).
 However, battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint.
-In this paper, we consider a composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery.
-This permits convex optimization of the composite battery SOC trajectory while ensuring admissibility of the resulting (aggregated) power schedule and disaggregation to the individual energy storage elements.
+In this paper, we develop a novel composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery.
+This permits convex optimization of the composite battery SOC trajectory while guaranteeing admissibility of the resulting (aggregated) power schedule and its disaggregation to the individual energy storage elements.
 Index Terms:
-Battery, energy storage, optimization, convex.
+Battery, energy storage, optimization, convex, priority-based control policy, complementarity conditions.
 I Introduction
-Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s transition to renewable energy and improve reliability  [Battery_review_paper].
-Grid-scale BESSs are comprised of several battery racks that can be split into several battery modules and further into battery cells.
-When optimizing BESS dispatch the individual elements are lumped as one composite battery and the composite dispatch is shared equally between the battery elements.
-Optimizing composite dispatch is a non-convex problem since a complementarity constraint is needed to avoid simultaneous charging and discharging.
-The complementarity constraint can be formulated with a non-linear constraint, or by the use of binary variables.
-To obtain a convex formulation, the complementarity constraint is typically omitted.
-The optimal power trajectories then exhibit simultaneous charging and discharging, but only the net power trajectory is applied.
-Implementing the net power trajectory results in a discrepancy between the predicted optimal SOC trajectory and the realized SOC trajectory [RobustNawaf, D.Pozo_paper].
-This discrepancy can lead to unexpected SOC saturation, resulting in un-realizable power trajectories.
-In [RobustNawaf, Mazen_CDC_paper], a robust linear battery model is presented as an inner approximation that guarantees that any optimal input power trajectory is realizable, but it is conservative over longer prediction horizons.
-Previous work has always assumed equal power-sharing between battery elements is enforced.
-This work revisits battery modeling by considering a composite battery made of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements.
-The paper characterizes the set of realizable composite power trajectories that can be disaggregated to the elements while ensuring element-wise power limits, energy limits, and complementarity constraints are satisfied.
-A priority-based strategy is employed to perform disaggregation.
+Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s uptake of variable renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation, BESS_renewables].
+In many applications, BESS dispatch involves coordinating $N$ battery racks/modules that together comprise a single BESS unit (e.g., DC-coupled elements), $N$ BESS container units in a utility-scale battery storage facility (e.g., AC-coupled elements), or $N$ aggregated BESS units coordinated as a virtual battery/power plant (VPP).
+However, when optimizing BESS dispatch, the individual elements that make up the BESS are often lumped together into a single composite energy storage asset.
+The composite dispatch is then optimized and the resulting power schedule is shared equally among the $N$ battery elements.
+An optimal composite dispatch represents a non-convex problem since the complementarity constraint is formulated with a non-linear constraint or by the use of binary variables [MILP_for_CC, prat_2024].
+To obtain a convex relaxation of the dispatch problem, the complementarity constraint is typically omitted [prat_2024].
+The resulting power trajectories can then exhibit simultaneous charging and discharging, which cannot be implemented.
+To recover near-term feasibility, only the net power trajectory is applied.
+As a result, the state of charge (SOC) becomes a provable lower bound on the SOC when the schedule is realized [RobustNawaf, almassalkhi_model-predictive_2015].
+This can lead to unexpected SOC saturation, resulting in an un-realizable power trajectory.
+Extensive research has investigated the conditions under which the convex relaxation’s state of charge (SOC) coincides with the actual realized SOC.
+Authors in [K.Baker_DCOPF_No_CC] avoid the use of the complementarity constraint when applied to DCOPF but add a penalty term to the objective function to ensure the exactness of the relaxed model.
+Ex-post sufficient conditions under which the relaxation becomes exact for an economic dispatch problem are provided in [Sufficient_Cond_for_Relax, Extended_Sufficient_Cond_for_Relax].
+Ex-ante conditions for exactness of the relaxed problem are presented in [Lin_Exact_Cond_Ex_Ante], assuming that certain constraints on battery power limits, as well as initial and maximum energy capacities, are met.
+Furthermore, [positive_prices, Jakob_Pos_Quadratic_Cost, prat_2024] shows that for economic dispatch problems with positive linear and quadratic cost functions, the convex relaxation at optimality would be exact with no simultaneous charging and discharging.
+However, the exactness of the relaxed model does not hold in general which may lead to battery saturation and inadmissible power trajectories.
+A simple example of this is the case with negative prices as shown in [Baldick_Neg_Price_Bad].
+For a single time step, this problem can be alleviated by developing tighter relaxations [D.Pozo_paper, elgersma2024], which can also heuristically be applied to time series, albeit without realizability guarantees.
+Figure 1: Comparing different composite models after implementing an open-loop predicted dispatch schedule.
+The relaxed model dispatch is inadmissible, while the robust linear model is admissible but suboptimal.
+However, this paper’s model achieves near-optimal performance and is admissible.
+Additionally, authors in  [RobustNawaf, Mazen_CDC_paper], an alternative robust linear battery model is presented which is a convex inner approximation of the non-convex formulation that sacrifices optimality for guaranteed admissibility.
+The robust linear model omits the complementarity constraint but envelopes the actual SOC trajectories to guarantee that any feasible power trajectory is realizable.
+However, the envelopes beget conservativeness as a function of efficiencies and prediction horizon, which limits practicality over longer horizons.
+In summary, previous work in energy storage optimization that circumvents the complementarity condition either has been application-specific and holds only if specific criteria on the objective function and/or the battery parameters are met, or presents convex restrictions that may yield significantly sub-optimal solutions.
+Furthermore, previous work has generally (and often, implicitly) assumed an equal power-sharing policy, which limits how individual elements can respond under dynamic conditions.
+Thus, this manuscript revisits the modeling of BESS aggregations by presenting a novel composite model $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements.
+The main contributions of the paper are as follows:
+  1. 1.
+    We present a novel linear model that provides a sufficient representation of aggregations of homogeneous battery storage elements.
+  2. 2.
+    We develop a priority-based strategy, based on element SOC, to perform the composite-to-individual disaggregation.
+  3. 3.
+    The proposed model guarantees that all individual element-wise power limits, energy limits, and complementarity constraints will be satisfied after disaggregation.
+Figure 1 illustrates the open-loop implementation of the dispatch schedules obtained from the proposed model, the relaxed model, and the robust linear model presented in [RobustNawaf] when used to optimize composite battery dispatch to track a given power reference signal.
+The relaxed model dispatch causes the battery to saturate leading to poor tracking performance.
+The robust linear model dispatch is realizable; however, yields sub-optimal tracking performance.
+The proposed model yields near-optimal results and guarantees realizable power trajectories under an unequal power-sharing policy.
+The rest of the paper is organized as follows.
+Section II presents the composite battery model and the associated constraints.
+Section III presents the proposed disaggregation strategy and the sufficient conditions for the admissibility of the composite power trajectory.
+Section IV presents simulation results to demonstrate the performance of the proposed model and disaggregation strategy.
+Finally, Section V concludes the paper.
 II Composite Battery Model
-We consider a generic BESS as shown in Fig 1.
+In this section, we will present element-wise and composite models of BESS systems.
+We consider BESS assets composed of an aggregated generic energy storage asset.
+An example of such a system is shown in Fig 2.
 It is composed of $N$ identical elements.
-Examples of this are DC-coupled battery racks connected to form a single BESS unit, or even multiple AC-coupled BESS units coordinated to form an aggregated (virtual) battery.
-Figure 1: Example of a composite battery with dc-coupled battery elements.
-For element $i$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$.
+Figure 2: Example of a composite battery with AC-coupled battery elements.
+For element $i\in\{1,\ldots,N\}$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$.
 The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOC at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively.
@@
 Note that all elements have identical power and energy limits.
-This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$.
-Battery element $i$
-$\begin{split}&E^{i}(l+1)=E^{i}(l)+\\ &\delta t\eta_{\text{c}}P_{\text{c}}^{i}(l)-\delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}(l)\end{split}$ (1a)
+This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$:
+$E^{i}(l+1)=E^{i}(l)+\delta t\eta_{\text{c}}P_{\text{c}}^{i}(l)-\delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}(l)$ (1a)
 $E^{i}(0)=E^{i}_{0}$ (1b)
-$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)$ (1c)
-$0\leq E^{i}(l+1)\leq{E}_{\text{max}}$ (1d)
-$0\leq P_{\text{c/d}}^{i}(l)\leq{P}_{\text{c/d,max}}$ (1e)
-Balanced composite battery
-$\begin{split}&E[k+1]=E[k]+\\ &\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]\end{split}$ (2a)
-$E[0]=\sum_{i}E^{i}_{0}$ (2b)
-$0=P_{\text{c}}[k]P_{\text{d}}[k]$ (2c)
-$0\leq E[k+1]\leq N{E}_{\text{max}}$ (2d)
-$0\leq P_{\text{c/d}}[k]\leq N{P}_{\text{c/d,max}}$ (2e)
+$0\leq E^{i}(l+1)\leq{E}_{\text{max}}$ (1c)
+$0\leq P_{\text{c/d}}^{i}(l)\leq{P}_{\text{c/d,max}}$ (1d)
+$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l).$ (1e)
 The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$.
@@
 $\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k].$ (2)
-Typically, a composite BESS is modeled and operated in a balanced way: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and equal power sharing across its elements.
-This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (2).
-Under equal power sharing, a single complementarity constraint at the composite battery level, (2c), is sufficient to satisfy (1c) for all $i$ and all $l$.
-Under balanced operation, a feasible composite dispatch implicitly satisfies all elements’ constraints in (1).
-However, under unequal power sharing, this does not hold.
-In fact, with unequal power sharing, we can omit the complementarity constraint at the composite level, but then have to disaggregate the composite power trajectory across the $N$ elements while satisfying (1c).
-That is, unequal power sharing increases the capability of the composite battery, but at the cost of computational complexity with $N$ complementary constraints.
-Confronting this challenge, this paper presents a priority-based disaggregation strategy that engenders a provably convex linear set of realizable composite power trajectories.
-These trajectories permit simultaneous charging and discharging of the BESS and approximate the full solution space arbitrarily closely in well-defined limits.
+Typically, it is assumed that a composite BESS system is modeled and operated under balanced conditions: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and adopts an equal power sharing policy across its elements.
+This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (3):
+$E[k+1]$ $=E[k]+\Delta t\eta_{\text{c}}P_{\text{c}}[k]-\Delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}[k]$ (3a)
+$E[0]$ $=\sum_{i}E^{i}_{0}$ (3b)
+00 0 $\leq E[k+1]\leq N{E}_{\text{max}}$ (3c)
+00 0 $\leq P_{\text{c/d}}[k]\leq N{P}_{\text{c/d,max}}$ (3d)
+00 0 $=P_{\text{c}}[k]P_{\text{d}}[k].$ (3e)
+Under the equal power sharing policy, a single complementarity constraint (3e) at the composite battery level is sufficient to satisfy (1e) for all $i$ and all $l$.
+Under balanced operation, a feasible composite dispatch based on (3) satisfies all elements’ constraints in (1).
+The complementarity constraint is often neglected to obtain a convex relaxation of the problem. For example, as
+[…]
+$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}$ $+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq 1,$ (4b)
+where the cutting plane in (4b) has been added to tighten the relaxation of the problem.
+However, such relaxations can result in an optimal power schedule with simultaneous charging and discharging, which cannot be implemented under equal load sharing.
+Tighter relaxations have been developed [D.Pozo_paper, elgersma2024] to reduce the prevalence of such solutions in real-world problems.
+Nevertheless, there are cases where simultaneous charging and discharging (which effectively ‘discards’ energy if $\eta_{c}\eta_{d}<1$) is a desirable outcome of a battery dispatch schedule.
+For example, this may be the case when energy prices are negative, when the storage unit tracks a dispatch signal, or as byproduct of minimum up- or downtime constraints in a centrally dispatched system.
+Such cases can be addressed by a composite battery if we omit the equal load sharing requirement.
+This way, we can realize simultaneous charging and discharging dispatch schedules by appropriately disaggregating them across the $N$ elements while satisfying (1e).
+This increases the capability of the composite battery at the cost of computational complexity with $N$ complementary constraints.
+In the remainder of this paper, we address this challenge by presenting a sufficient (but not unique) priority-based disaggregation strategy that engenders a linear set of provably realizable composite power trajectories.
+These trajectories permit simultaneous charging and discharging of the BESS and approximate the full solution space arbitrarily closely.
 III Composite Trajectory Disaggregation
@@
 III-A Priority Stack Controller (PSC)
-Every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC.
+For every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC.
 The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority.
@@
 III-B Guaranteeing Element Complementarity Constraints
-Next, we analyze the effects of the PSC on (1), on complementarity conditions (1c), and on energy limits (1d).
+Next, we analyze the effects of the PSC on (1), on complementarity conditions (1e), and on energy limits (1c).
 To ensure the complementarity constraints are satisfied at the element level, $(P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l)=0\quad\forall i,\forall l)$, the priority stack cannot issue a charging and discharging demand to the same battery element $i$ simultaneously.
-Thus, any overlap in the selection of $N_{\text{c}}(l)$ charging and $N_{\text{d}}(l)$ discharging elements imply a violation of the complementarity constraints, which must, therefore, satisfy:
-$N_{\text{c}}(l)+N_{\text{d}}(l)\leq N\quad\forall l.$ (5)
+Thus, any overlap in the selection of $N_{\text{c}}(l)$ charging and $N_{\text{d}}(l)$ discharging elements imply a violation of the complementarity constraints, which must, therefore, satisfy
+$N_{\text{c}}(l)+N_{\text{d}}(l)\leq N,\quad\forall l.$ (6)
 Proposition 1.
 Under the PSC, if $P_{\text{c}}[k]$ and $P_{\text{d}}[k]$ satisfy
-$\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\leq N-1.$ (6)
+$\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\leq N-1,$ (7)
 then the resulting $N_{\text{c}}(l)$ and $N_{\text{d}}(l)$ satisfy (6).
@@
 $\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}\leq\Big{\lceil}N-1-(\Gamma_{\text{c}}+\Delta_{\text{c}})\Big{\rceil}=N-1-\Gamma_{\text{c}}.$ (8)
-Adding inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$ for any $l$. ∎
+Adding the inequalities, we get $N_{c}(l)+N_{d}(l)\leq N$ $\forall l$. ∎
 Thus, adding (7) to the composite formulation in (3) will satisfy element complementarity constraints under PSC. The feasible regions dictated by (6) and (7) are shown in Fig. 3.
-Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are not considered here.
+Figure 3: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are omitted here.
 III-C Relating Element and Composite Energy Limit Constraints
@@
 with
-$\varepsilon:=\frac{\Delta t}{M}\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right),$ (9)
-will not violate any individual element’s energy limit, provided that $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$.
+$\varepsilon:=\delta t\left(\eta_{\text{c}}{P}_{\text{c,max}}+\frac{{P}_{\text{d,max}}}{\eta_{\text{d}}}\right),$ (10)
+will not violate any individual element’s energy limit, provided that (i) $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$ and (ii) the control time step $\delta t=\Delta t/M$ is sufficiently small to ensure $\varepsilon\leq\frac{1}{2}E_{\text{max}}$.
 Proof.
@@
 $\frac{\Delta t}{M}\eta_{\text{c}}\left(P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\right)$ $-\frac{\Delta t}{M\eta_{\text{d}}}\left(P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\right).$
-From Property 1 and (1e), we have $-P_{\text{c,max}}\leq P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\leq 0$ and $0\leq P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\leq P_{\text{d,max}}$, which implies
+From Property 1 and (1d), we have $-P_{\text{c,max}}\leq P^{i}_{\text{c}}(l)-P^{j}_{\text{c}}(l)\leq 0$ and $0\leq P^{i}_{\text{d}}(l)-P^{j}_{\text{d}}(l)\leq P_{\text{d,max}}$, which implies
 $E^{i}(l+1)-E^{j}(l+1)\leq$ $E^{i}(l)-E^{j}(l)\leq\Delta E(l),$ (12a)
@@
 Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$.
-Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$, we maintain $\Delta E(l)\leq\varepsilon$ for all $l$.
+Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$ (condition (i)), we maintain $\Delta E(l)\leq\varepsilon,\forall l$.
 Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (9) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$.
@@
 Because the SOC spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$.
-Combined, these inequalities show that (1d) holds for all $i$ and all $l$.
+Combined, these inequalities show that (1c) holds for all $i$ and all $l$.
+Lastly, we require condition (ii) ($\varepsilon\leq\frac{1}{2}E_{\text{max}}$) to ensure that (9) does not result in an empty set (infeasibility).
+This can is achieved with a sufficiently large value of $M$.
 ∎
@@
 […]
-$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}$ $\leq\frac{N-1}{N}$ $\forall k=0,\ldots,K-1$ (12b)
-$N\varepsilon\leq E[k]$ $\leq N\left(E_{\text{max}}-\varepsilon\right)$ $\forall k=0,\ldots,K$ (12c)
-$P_{\text{c}}[k],P_{\text{d}}[k]$ $\geq 0$ $\forall k=0,\ldots,K-1,$ (12d)
+[…]
+$\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq\frac{N-1}{N},$ $k=0,\ldots,K-1$ (13c)
+$N\varepsilon\leq E[k]\leq N\left(E_{\text{max}}-\varepsilon\right),$ $k=0,\ldots,K$ (13d)
+$P_{\text{c}}[k],P_{\text{d}}[k]\geq 0,$ $k=0,\ldots,K-1$ (13e)
 will satisfy (1) for all $N$ elements under PSC and is, thus, realizable.
@@
 (1a) and (2) imply (3a) and (1b) implies (3b).
-From Proposition 1, (12b) ensures satisfaction of (1c).
-Similarly, with $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, Proposition 2 guarantees that (1d) holds, and the PSC with (12d) and (3) ensure that (1e) holds for all $N$ elements.
-Thus, any feasible composite dispatch is realizable.∎
-Note that for a fixed composite battery, if $N$ increases, the cutting plane in (12b) approaches the cutting plane of the relaxed model.
+From Proposition 1, (13c) ensures satisfaction of (1e).
+Similarly, with $\max_{i,j}\{|E^{i}_{0}-E^{j}_{0}|\}\leq\varepsilon$, Proposition 2 guarantees that (1c) holds, and the PSC with (13e) and (2) ensure that (1d) holds for all $N$ elements.
+Thus, any feasible composite dispatch is realizable.
+∎
+Note that for a fixed composite battery, if $N$ increases, the cutting plane in (13c) approaches the cutting plane in (4b).
 Furthermore, for $M\gg 1$, $\varepsilon$ in (10) tends to zero.
-That is, with faster control and more granular elements, (12) approaches the relaxed battery model defined in [Mazen_CDC_paper]; however, with guaranteed realizability.
+That is, with faster control and more granular elements, (13) approaches the relaxed battery model defined in (4); however, with guaranteed realizability.
+The energy buffer $\varepsilon$ in (13d) restricts the set of feasible energy trajectories relative to (3) and (4a).
+The feasible energy trajectories in the robust linear model presented in [RobustNawaf] are also a subset of the feasible energy trajectories in (3) and (4a).
+However, the proposed model in (13) has a constant energy buffer $\varepsilon$ that is independent of the length of time horizon considered.
+This is in contrast to the robust model which has an energy buffer that increases with the length of the time horizon which, in the worst case, significantly reduces the feasible set of energy trajectories and can yield sub-optimal solutions.
+This is illustrated in Fig.
+4 where the (worst case) feasible set of energy trajectories for the robust model and the RCB model for $\Delta t=1\text{hr}$ and $M=1$ ($\delta t=1\text{hr}$) and $M=10$ ($\delta t=6\text{min}$) are compared for the battery parameters of a Tesla Powerwall shown in Table I.Clearly, smaller control time steps result in a smaller $\varepsilon$ and thus a larger feasible set of energy trajectories for the proposed model, while the robust model remains unchanged.
+Finally, we note that the PSC policy is a greedy sufficient policy that fully activates BESS elements in sequence.
+More elaborate policies (greedy or not) can be developed that preferentially spread element loading for most aggregate dispatch signals, without violating the conditions of Theorem 1.
+TABLE I: Individual BESS Element Parameters
+Parameters Value Unit
+Charge/discharge efficiencies, $\eta_{\text{c}}/\eta_{\text{d}}$ 0.95 -
+Maximum charge/discharge power, $P_{\text{max}}$ 5.00 kW
+Maximum energy capacity, $E_{\text{max}}$ 13.5 kWh
+Figure 4: Feasible set energy trajectories for the relaxed model, the robust model and the proposed realizable composite battery model (RCB).
 IV Simulation Results
-We apply the realizable composite battery model (RCB) in (12) to minimize power tracking error for a given reference signal based on the objective function $\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}$.
-For comparison, relaxed and robust models from [RobustNawaf] are used under balanced operation.
+We apply the realizable composite battery (RCB) model in (13) to two relevant use cases and compare its performance against the tighter relaxed model given in (4) and the robust model in [RobustNawaf] used under balanced operation.
 Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison.
-The chosen element battery parameters are obtained from a Tesla Powerwall and we set $N=1000$ and $M=1$.
+The battery element parameters given in Table I were used.
+The two use-cases are a) power reference tracking and b) revenue maximization.
+IV-A Power Reference Tracking
+In this use case, the goal is to minimize the power tracking error for a given reference signal as shown below,
+$\min_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}.$ (14)
+A timestep of $\Delta t=5$ min, along with $N=100$ was used.
+The RCB model was used three times with $M=1$, $M=5$, and $M=10$ respectively.
 The results from using each model are shown in Fig.
-3.
-Figure 3: Reference tracking optimization.
-The RCB model and relaxed battery model produced similar optimal dispatch.
-However, this optimal dispatch was not realizable under equal power sharing (relaxed) causing the battery to saturate, resulting in the sudden drop in power seen in Fig.
-3.
-The MILP, robust, and RCB models produced realizable results, with the robust formulation resulting in worse tracking performance.
-The MILP and RCB formulations produced comparable results, however, with a significant computational advantage for the RCB model.
+5 and the solve times along with the mean squared tracking error are shown in Table II.
+Figure 5: Open-loop implementation of power optimal reference tracking power schedules using MILP, relaxed, robust and proposed RCB models.
+TABLE II: Power Reference Tracking Performance Metrics
+Model Solver Predicted Actual
+Time (ms) MSE (kW2) MSE (kW2)
+MILP 50,140 0 0
+Relaxed 30 0 1841
+Robust 34 2.12 2.12
+RCB (M = 1) 30 0.72 0.72
+RCB (M = 5) 30 0.013 0.013
+RCB (M = 10) 30 0 0
+The MILP model yielded the optimal solution and was able to track the reference signal with no tracking error.
+However, the MILP also came with a significant computational burden compared to the rest of the formulations.
+The other formulations are linear and, as a result, have significantly faster solve times.
+The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs leading to a large realized tracking error.
+The robust model yields a realizable yet sub-optimal solution.
+The RCB model shows improved tracking error with increasing values of $M$ but even with $M=1$ the tracking error surpasses the robust model.
+IV-B Revenue Maximization
+In this use case, the goal is to maximize the predicted revenue from the composite battery system given a forecasted price signal. The objective function is given by,
+$\max_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K}C[k](P_{\text{d}}[k]-P_{\text{c}}[k]),$ (15)
+where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices for 24 hours, plotted in Fig 6.
+Figure 6: CAISO day-ahead price forecasts
+A timestep of $\Delta t=15$ min, along with $N=100$ was used.
+The RCB model was tested for $M=$ 1, 5, and 10 which correspond to $\delta t=$ 15 min, 3 min and 1.5 min respectively.
+An additional case was added with $M=900$ ($\delta t=1s$) to gain insight on model performance as $M\rightarrow\infty.$The results for the different models are shown in Table III.
+TABLE III: Revenue Maximization Performance metrics
+Model Solver Predicted Actual
+Time (ms) Revenue (\$) Revenue (\$)
+MILP 1250 2088.97 2088.97
+Relaxed 1 2092.80 1940.35
+Robust 1 1866.42 1866.42
+RCB (M = 1) 1 1411.53 1411.53
+RCB (M = 5) 1 1955.76 1955.76
+RCB (M = 10) 1 2022.87 2022.87
+RCB (M = 900) 1 2089.23 2089.23
+Similar to the previous use-case, the MILP model produced the highest revenue but at a significantly larger computational cost.
+The relaxed model predicts the highest revenue but underestimates the SOC, leading to unpredicted battery saturation and a realized performance that is significantly worse.
+The robust model yields a realizable yet suboptimal solution.
+The RCB model for small values of $M$ results in a sub-optimal solution which can be attributed to the larger timestep used in this use-case leading to a larger value or $\varepsilon$ and, in turn, a more restricted feasible set.
+However, for f $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and approaches the predicted performance for larger values of $M$.
+It was also observed that the RCB surpasses the robust model for $M\geq 3$ ($\delta t\leq 5$ min).
 V Conclusion
@@
 By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories.
-Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems.
+Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems or aggregate storage representations in energy system optimization models.
 […]

### v6

@@
-Linear Model of Aggregated Homogeneous Energy Storage Elements with Realizable Dispatch Guarantees
+Linear Aggregate Model for Realizable Dispatch of Homogeneous Energy Storage
 Abstract
-To optimize the dispatch of batteries, a model is required that can predict the state of charge (SOC) trajectory for a chosen open-loop power schedule to ensure admissibility (i.e., that schedule can be realized).
-However, battery dispatch optimization is inherently challenging since batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint.
+To optimize the dispatch of batteries, a model is required that can predict the state of energy (SOE) trajectory for a chosen open-loop power schedule to ensure admissibility (i.e., that schedule can be realized).
+However, battery dispatch optimization is inherently challenging when batteries cannot simultaneously charge and discharge, which begets a non-convex complementarity constraint.
 In this paper, we develop a novel composition of energy storage elements that can charge or discharge independently and provide a sufficient linear energy storage model of the composite battery.
-This permits convex optimization of the composite battery SOC trajectory while guaranteeing admissibility of the resulting (aggregated) power schedule and its disaggregation to the individual energy storage elements.
+This permits convex optimization of the composite battery dispatch while ensuring the admissibility of the resulting (aggregated) power schedule and its disaggregation to the individual elements.
 Index Terms:
-Battery, energy storage, optimization, convex, priority-based control policy, complementarity conditions.
+Battery, energy storage, convex optimization, priority-based control, complementarity constraints.
 I Introduction
 Battery energy storage systems (BESS) are increasingly deployed to help facilitate the grid’s uptake of variable renewable energy and improve reliability  [Battery_review_paper, BESS_simultaneous_participation, BESS_renewables].
-In many applications, BESS dispatch involves coordinating $N$ battery racks/modules that together comprise a single BESS unit (e.g., DC-coupled elements), $N$ BESS container units in a utility-scale battery storage facility (e.g., AC-coupled elements), or $N$ aggregated BESS units coordinated as a virtual battery/power plant (VPP).
-However, when optimizing BESS dispatch, the individual elements that make up the BESS are often lumped together into a single composite energy storage asset.
-The composite dispatch is then optimized and the resulting power schedule is shared equally among the $N$ battery elements.
-An optimal composite dispatch represents a non-convex problem since the complementarity constraint is formulated with a non-linear constraint or by the use of binary variables [MILP_for_CC, prat_2024].
-To obtain a convex relaxation of the dispatch problem, the complementarity constraint is typically omitted [prat_2024].
-The resulting power trajectories can then exhibit simultaneous charging and discharging, which cannot be implemented.
-To recover near-term feasibility, only the net power trajectory is applied.
-As a result, the state of charge (SOC) becomes a provable lower bound on the SOC when the schedule is realized [RobustNawaf, almassalkhi_model-predictive_2015].
-This can lead to unexpected SOC saturation, resulting in an un-realizable power trajectory.
-Extensive research has investigated the conditions under which the convex relaxation’s state of charge (SOC) coincides with the actual realized SOC.
-Authors in [K.Baker_DCOPF_No_CC] avoid the use of the complementarity constraint when applied to DCOPF but add a penalty term to the objective function to ensure the exactness of the relaxed model.
-Ex-post sufficient conditions under which the relaxation becomes exact for an economic dispatch problem are provided in [Sufficient_Cond_for_Relax, Extended_Sufficient_Cond_for_Relax].
-Ex-ante conditions for exactness of the relaxed problem are presented in [Lin_Exact_Cond_Ex_Ante], assuming that certain constraints on battery power limits, as well as initial and maximum energy capacities, are met.
-Furthermore, [positive_prices, Jakob_Pos_Quadratic_Cost, prat_2024] shows that for economic dispatch problems with positive linear and quadratic cost functions, the convex relaxation at optimality would be exact with no simultaneous charging and discharging.
-However, the exactness of the relaxed model does not hold in general which may lead to battery saturation and inadmissible power trajectories.
-A simple example of this is the case with negative prices as shown in [Baldick_Neg_Price_Bad].
-For a single time step, this problem can be alleviated by developing tighter relaxations [D.Pozo_paper, elgersma2024], which can also heuristically be applied to time series, albeit without realizability guarantees.
-Figure 1: Comparing different composite models after implementing an open-loop predicted dispatch schedule.
-The relaxed model dispatch is inadmissible, while the robust linear model is admissible but suboptimal.
-However, this paper’s model achieves near-optimal performance and is admissible.
-Additionally, authors in  [RobustNawaf, Mazen_CDC_paper], an alternative robust linear battery model is presented which is a convex inner approximation of the non-convex formulation that sacrifices optimality for guaranteed admissibility.
-The robust linear model omits the complementarity constraint but envelopes the actual SOC trajectories to guarantee that any feasible power trajectory is realizable.
-However, the envelopes beget conservativeness as a function of efficiencies and prediction horizon, which limits practicality over longer horizons.
+In many applications, BESS dispatch involves coordinating large numbers of energy storage elements.
+Examples include battery racks that comprise a single BESS asset, BESS container assets in a utility-scale battery storage facility as shown in Fig 1, or aggregated BESS units coordinated as a virtual power plant (VPP).
+Optimally scheduling the charging and discharging of storage elements requires solving a non-convex optimization problem.
+Because most storage technologies do not permit simultaneous charging and discharging, complementarity constraints are usually required to force the product of the charging and discharging powers to be equal to zero at each time step, and for each element.
+This nonlinearity is often implemented using binary variables [MILP_for_CC, prat_2024].
+The non-convexity of the composite BESS dispatch can present a computational challenge when incorporating BESSs into more complex, large-scale power system optimization problems, such as optimal power flow (OPF) and economic dispatch [qi2025locationalenergystoragebid].
+Figure 1: Example of a composite battery with AC-coupled battery elements.
+This motivates the need for a convex BESS model.
+A common approach is to obtain a convex relaxation by omitting the complementarity constraint, which results in a predicted state of energy (SOE) that is a provable lower bound on the realized SOE and can lead to unexpected SOE saturation and an inadmissible dispatch [RobustNawaf, almassalkhi_model-predictive_2015].
+Additionally, the relaxation permits solutions that simultaneously charge and discharge [prat_2024].
+At first glance, such solutions may appear irrational because charging/discharging inefficiencies result in unnecessary dissipation of energy.
+Nevertheless, there are cases where such solutions are a rational outcome of a battery dispatch schedule, e.g., when energy prices are negative [Baldick_Neg_Price_Bad], or as a byproduct of minimum up- or downtime constraints in a centrally dispatched system.
+Any dissipative effects on battery aging can be captured by adding a cycling cost to the objective function, so that the benefit must outweigh the cost.
+Sufficient conditions under which simultaneously charging and discharging is optimal (i.e., rational) are given in [prat_2024].
+Conversely, there are cases where simultaneous charging and discharging is guaranteed not to occur.
+Sufficient ex-post conditions for an exact relaxation in economic dispatch are provided in [Sufficient_Cond_for_Relax].
+Ex-ante conditions for exactness of the relaxed problem are presented in [Lin_Exact_Cond_Ex_Ante], assuming that battery power limits, as well as initial and maximum energy capacities, satisfy certain conditions.
+Furthermore, [positive_prices, Jakob_Pos_Quadratic_Cost, prat_2024] show that for specific economic dispatch formulations with positive linear or quadratic cost functions, the relaxation at optimality would be exact with no simultaneous charging and discharging.
+Various alternative problem formulations exist to reduce the reliance on the computationally demanding complementarity constraint.
+For a single time step, tighter relaxations have been proposed [D.Pozo_paper, elgersma2024], which can also heuristically be applied to time series, albeit without realizability guarantees.
+Authors in [K.Baker_DCOPF_No_CC] avoid the use of the complementarity constraint when applied to DC Optimal Power Flow (DCOPF) but add a penalty term to the objective function to ensure the exactness of the relaxed model.
+Instead of modifying objective, the authors in  [RobustNawaf, Mazen_CDC_paper] present an alternative robust linear battery model which represents a convex restriction of the non-convex formulation that sacrifices optimality for guaranteed admissibility.
+The robust linear model omits the complementarity constraint but envelopes the actual SOE trajectories to guarantee that any feasible power trajectory is realizable.
+However, the envelopes beget conservativeness as a function of energy conversion efficiencies and the prediction horizon, which limits practicality.Now, a composite battery made up of identical storage elements can be scheduled with various methods.
+One way is to model and optimize the schedule of each storage element individually, requiring many complementarity constraints [equal_unequal].
+A simpler and computationally efficient approach is to use a static equal power sharing policy where all elements charge or discharge equally [equal_unequal, equal_unequal2].
+The elements can then be aggregated into a composite battery whose dispatch is optimized, but still requires one non-convex complementarity constraint.
 In summary, previous work in energy storage optimization that circumvents the complementarity condition either has been application-specific and holds only if specific criteria on the objective function and/or the battery parameters are met, or presents convex restrictions that may yield significantly sub-optimal solutions.
-Furthermore, previous work has generally (and often, implicitly) assumed an equal power-sharing policy, which limits how individual elements can respond under dynamic conditions.
-Thus, this manuscript revisits the modeling of BESS aggregations by presenting a novel composite model $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements.
+Furthermore, previous work has generally (and often, implicitly) assumed an equal power-sharing policy, which limits how elements can be dispatched.
+This manuscript revisits the modeling of aggregations of storage units by presenting a novel composite model of $N$ identical and independently controllable battery elements, where simultaneous charging and discharging power trajectories at the composite level can be implemented by inhomogeneous disaggregation to the individual elements.
+Throughout, BESS aggregations are used as an example, but the proposed work can be extended to different energy storage assets.
 The main contributions of the paper are as follows:
   1. 1.
-    We present a novel linear model that provides a sufficient representation of aggregations of homogeneous battery storage elements.
+    We present a novel linear model that provides a sufficient representation of aggregations of homogeneous energy storage elements.
+For large numbers of elements or and small control time steps, the representation asymptotically approaches the true feasibility envelope.
   2. 2.
-    We develop a priority-based strategy, based on element SOC, to perform the composite-to-individual disaggregation.
+    We develop a priority-based strategy, based on element SOE, to perform composite-to-elements disaggregation.
   3. 3.
-    The proposed model guarantees that all individual element-wise power limits, energy limits, and complementarity constraints will be satisfied after disaggregation.
-Figure 1 illustrates the open-loop implementation of the dispatch schedules obtained from the proposed model, the relaxed model, and the robust linear model presented in [RobustNawaf] when used to optimize composite battery dispatch to track a given power reference signal.
-The relaxed model dispatch causes the battery to saturate leading to poor tracking performance.
-The robust linear model dispatch is realizable; however, yields sub-optimal tracking performance.
-The proposed model yields near-optimal results and guarantees realizable power trajectories under an unequal power-sharing policy.
-The rest of the paper is organized as follows.
-Section II presents the composite battery model and the associated constraints.
-Section III presents the proposed disaggregation strategy and the sufficient conditions for the admissibility of the composite power trajectory.
-Section IV presents simulation results to demonstrate the performance of the proposed model and disaggregation strategy.
-Finally, Section V concludes the paper.
+    We prove that all solutions of the linear aggregate model are realizable using the priority-based strategy, respecting element-wise power limits, energy limits, and complementarity constraints after disaggregation.
+Next, we develop the composite battery model in Section II.
+Section III presents the proposed disaggregation strategy and sufficient conditions for the admissibility of a composite power trajectory.
+Simulation results that demonstrate the performance of the proposed model and disaggregation strategy make up Section IV.
+Conclusion and future work are in Section V.
 II Composite Battery Model
-In this section, we will present element-wise and composite models of BESS systems.
-We consider BESS assets composed of an aggregated generic energy storage asset.
-An example of such a system is shown in Fig 2.
-It is composed of $N$ identical elements.
-Figure 2: Example of a composite battery with AC-coupled battery elements.
+In this section, we will present element-wise and composite models of BESS systems, composed of $N$ identical and independently controllable elements (e.g., as shown in Fig 1).
 For element $i\in\{1,\ldots,N\}$, the charge/discharge power is set in controller time steps $l$ with duration $\delta t$.
-The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOC at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively.
+The charge/discharge power between $l$ and $l+1$, the charge/discharge efficiency and the SOE at timestep $l$ are denoted by $P^{i}_{\text{c/d}}(l)$, $\eta_{\text{c/d}}$, and $E^{i}(l)$, respectively.
 A realizable BESS dispatch must satisfy power limits (${P}_{\text{c/d,max}}$), energy limits (${E}_{\text{max}}$), and complementarity constraints for all battery elements ($\forall i=1,\ldots,N$) and time steps ($\forall l=0,\ldots,L-1$).
 Note that all elements have identical power and energy limits.
-This results in (1), where $E^{i}_{0}$ is the initial SOC of battery element $i$:
+This results in (1), where $E^{i}_{0}$ is the initial SOE of battery element $i$:
 $E^{i}(l+1)=E^{i}(l)+\delta t\eta_{\text{c}}P_{\text{c}}^{i}(l)-\delta t\frac{1}{\eta_{\text{d}}}P_{\text{d}}^{i}(l)$ (1a)
@@
 $0\leq P_{\text{c/d}}^{i}(l)\leq{P}_{\text{c/d,max}}$ (1d)
-$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l).$ (1e)
+$0=P_{\text{c}}^{i}(l)P_{\text{d}}^{i}(l),$ (1e)
+where (1c)-(1e) represent energy limits, power limits and, complementarity constraints, respectively.
 The scheduling of the composite battery is done with (optionally) longer scheduler time steps of duration $\Delta t=M\delta t,M\in\mathbb{N}^{+}$.
@@
 $\sum_{i=1}^{N}P_{\text{c}}^{i}(l)=P_{\text{c}}[k]\quad\land\quad\sum_{i=1}^{N}P_{\text{d}}^{i}(l)=P_{\text{d}}[k].$ (2)
-Typically, it is assumed that a composite BESS system is modeled and operated under balanced conditions: the dispatcher (e.g., EMS) maintains equal SOC ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and adopts an equal power sharing policy across its elements.
+Typically, it is assumed that a composite BESS system is modeled and operated under balanced conditions: the dispatcher (e.g., EMS) maintains equal SOE ($E_{0}^{i}=E_{0}~{}~{}\forall i$) and adopts an equal power sharing policy across its elements.
 This results in a composite battery with power and energy limits of $N{P}_{\text{c/d,max}}$ and $N{E}_{\text{max}}$ as shown in (3):
@@
 00 0 $=P_{\text{c}}[k]P_{\text{d}}[k].$ (3e)
-Under the equal power sharing policy, a single complementarity constraint (3e) at the composite battery level is sufficient to satisfy (1e) for all $i$ and all $l$.
-Under balanced operation, a feasible composite dispatch based on (3) satisfies all elements’ constraints in (1).
+Under equal power sharing, a complementarity constraint (3e) at the composite battery level is sufficient to satisfy (1e) for all $i$ and all $l$.
+Under balanced operation, a composite dispatch satisfying (3) would satisfy (1) for all $i$ and all $l$..
 The complementarity constraint is often neglected to obtain a convex relaxation of the problem. For example, as
@@
 $\frac{P_{\text{c}}[k]}{N{P}_{\text{c,max}}}$ $+\frac{P_{\text{d}}[k]}{N{P}_{\text{d,max}}}\leq 1,$ (4b)
-where the cutting plane in (4b) has been added to tighten the relaxation of the problem.
+where the cutting plane in (4b) is one of the facets of the convex hull of the feasible set defined by (3), typically used to tighten the relaxation.
+Even tighter relaxations have been developed [D.Pozo_paper, elgersma2024] to reduce the prevalence of solutions with simultaneous charging and discharging in real-world problems.
 However, such relaxations can result in an optimal power schedule with simultaneous charging and discharging, which cannot be implemented under equal load sharing.
-Tighter relaxations have been developed [D.Pozo_paper, elgersma2024] to reduce the prevalence of such solutions in real-world problems.
-Nevertheless, there are cases where simultaneous charging and discharging (which effectively ‘discards’ energy if $\eta_{c}\eta_{d}<1$) is a desirable outcome of a battery dispatch schedule.
-For example, this may be the case when energy prices are negative, when the storage unit tracks a dispatch signal, or as byproduct of minimum up- or downtime constraints in a centrally dispatched system.
 Such cases can be addressed by a composite battery if we omit the equal load sharing requirement.
@@
 III-A Priority Stack Controller (PSC)
-For every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOC.
-The element with the lowest SOC has the highest charging priority while the element with the highest SOC has the highest discharging priority.
-When selecting the highest-priority element for charging/discharging that element’s power is maximized, i.e., $P^{i}_{\text{c/d}}(l)=P_{\text{c/d,max}}$, before actuating another charging/discharging element from the stack.
-This process is repeated every time-step $l$, adapting to changing element SOCs.
-The priority stack results in $N_{\text{c}}$ elements with lowest SOCs charging and $N_{\text{d}}$ elements with highest SOCs discharging, as follows:
+For every controller time-step $l$, the priority stack controller (PSC) sorts all $N$ elements based on the SOE.
+The element with the lowest SOE has the highest charging priority, while the element with the highest SOE has the highest discharging priority.
+Note that similar control schemes have been proposed in [karan_kalsi] and investigated in a real-time operational setting in [equal_unequal, equal_unequal2].
+The priority stack results in $N_{\text{c}}$ elements with lowest SOEs charging and $N_{\text{d}}$ elements with highest SOEs discharging, as follows:
 $N_{\text{c}}(l):=\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\quad\land\quad N_{\text{d}}(l):=\Big{\lceil}\frac{P_{\text{d}}[k]}{{P}_{\text{d,max}}}\Big{\rceil}.$ (5)
+Disaggregating the composite charging trajectory for $P_{\text{c}}[k]>0$, the PSC sets $P^{i}_{\text{c}}(l)=P_{\text{c,max}}$ for the $N_{\text{c}}(l)-1$ elements with the lowest SOE and the next-lowest element is assigned $P^{i}_{\text{c}}(l)=P_{\text{c}}[k]-(N_{\text{c}}(l)-1)P_{\text{c,max}}$.
+Discharging is implemented analogously.
 The PSC is summarized in Algorithm 1 and begets Property 1.
@@
 5:    Assign $P^{i}_{\text{c/d}}(l)$ for first $N_{\text{c}}(l)$/last $N_{\text{d}}(l)$ elements
-6:    Update SOC for all elements: $E^{i}(l)\rightarrow E^{i}(l+1)$
+6:    Update SOE for all elements: $E^{i}(l)\rightarrow E^{i}(l+1)$
 7:end for
@@
 Proof.
-Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$.
+Define $\Gamma_{\text{c}}:=\lfloor\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\rfloor$ and $\Delta_{\text{c}}:=\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}-\Gamma_{\text{c}}\in[0,1)$ to be the integer and fractional components of the quantity $\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}$.
 Then $\Big{\lceil}\frac{P_{\text{c}}[k]}{{P}_{\text{c,max}}}\Big{\rceil}\leq\Gamma_{c}+1$ and, from (7), we get
@@
 Thus, adding (7) to the composite formulation in (3) will satisfy element complementarity constraints under PSC. The feasible regions dictated by (6) and (7) are shown in Fig. 2.
-Figure 3: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOC limits are omitted here.
-III-C Relating Element and Composite Energy Limit Constraints
+Figure 2: Feasible region for $N=10$ and $P_{\text{c/d,max}}=0.5$MW (left) and $N=30$ and $P_{\text{c/d,max}}=0.17$MW (right). SOE limits are omitted here.
+III-C Relating Element and Composite Energy Constraints
 Similarly, we want to relate element energy limits to those in the composite formulation.
-Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOC trajectories.
-The spread of element SOCs can impact realizability of the composite dispatch when individual elements hit their SOC limits.
-Constraint violations can be avoided by quantifying the maximum possible SOC spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
+Under PSC, the battery elements are charged/discharged individually every $l$, which results in differing SOE trajectories.
+The spread of element SOEs can impact the realizability of the composite dispatch when individual elements hit their SOE limits.
+Constraint violations can be avoided by quantifying the maximum possible SOE spread under PSC and using that as an energy buffer applied to the composite battery energy limits.
 Proposition 2.
@@
 Proof.
-Define the maximum difference in SOC across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOC spread).
-Consider element SOCs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$.
+Define the maximum difference in SOE across all elements as $\Delta E(l):=\max_{i,j}\{|E^{i}(l)-E^{j}(l)|\}$ (i.e., the SOE spread).
+Consider element SOEs $E^{i}(l)$ and $E^{j}(l)$ and, without loss of generality, let $E^{i}(l)\geq E^{j}(l)$.
 Using (1a) the following is obtained,
@@
 Considering all $i,j$, we can bound $\Delta E(l+1)\leq\max\left\{\Delta E(l),\varepsilon\right\}$.
-Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOC spread $\Delta E(0)\leq\varepsilon$ (condition (i)), we maintain $\Delta E(l)\leq\varepsilon,\forall l$.
+Moreover, we see that $\Delta E(l)\leq\varepsilon\Rightarrow\Delta E(l+1)\leq\varepsilon$, so it follows that for a composite battery with initial SOE spread $\Delta E(0)\leq\varepsilon$ (condition (i)), we maintain $\Delta E(l)\leq\varepsilon,\forall l$.
 Because energy linearly increases/decreases within each scheduling time step $k$ under PSC, (9) implies its equivalent for controller time steps $l$: $N\varepsilon\leq E(l)\leq N(E_{\text{max}}-\varepsilon)$.
-Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOC.
-Because the SOC spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$.
-Combined, these inequalities show that (1c) holds for all $i$ and all $l$.
+Dividing by $N$ begets $\varepsilon\leq\overline{E}(l)\leq E_{\text{max}}-\varepsilon$, where $\overline{E}(l):=\sum_{j}E^{j}(l)/N$ is the average element SOE.
+Because the SOE spread $\Delta E(l)\leq\varepsilon$, we must have $\overline{E}(l)-\varepsilon\leq E^{i}(l)\leq\overline{E}(l)+\varepsilon,\forall i$.
+Thus, (1c) holds for all $i$ and all $l$.
 Lastly, we require condition (ii) ($\varepsilon\leq\frac{1}{2}E_{\text{max}}$) to ensure that (9) does not result in an empty set (infeasibility).
@@
 Furthermore, for $M\gg 1$, $\varepsilon$ in (10) tends to zero.
-That is, with faster control and more granular elements, (13) approaches the relaxed battery model defined in (4); however, with guaranteed realizability.
+That is, with faster control and more granular elements, (13) approaches the relaxed model in (4); however, with guaranteed realizability.
 The energy buffer $\varepsilon$ in (13d) restricts the set of feasible energy trajectories relative to (3) and (4a).
@@
 This is illustrated in Fig.
-4 where the (worst case) feasible set of energy trajectories for the robust model and the RCB model for $\Delta t=1\text{hr}$ and $M=1$ ($\delta t=1\text{hr}$) and $M=10$ ($\delta t=6\text{min}$) are compared for the battery parameters of a Tesla Powerwall shown in Table I.Clearly, smaller control time steps result in a smaller $\varepsilon$ and thus a larger feasible set of energy trajectories for the proposed model, while the robust model remains unchanged.
+3 where the (worst case) feasible set of energy trajectories for the robust model and the RCB model for $\Delta t=1\text{hr}$ and $M=1$ ($\delta t=1\text{hr}$) and $M=5$ ($\delta t=12\text{min}$) are compared for the battery parameters of a Tesla Powerwall shown in Table I.
+Clearly, smaller control time steps result in a smaller $\varepsilon$ and thus a larger feasible set of energy trajectories for the proposed model, while the robust model remains unchanged.
 Finally, we note that the PSC policy is a greedy sufficient policy that fully activates BESS elements in sequence.
@@
 We apply the realizable composite battery (RCB) model in (13) to two relevant use cases and compare its performance against the tighter relaxed model given in (4) and the robust model in [RobustNawaf] used under balanced operation.
-Also, a Mixed-Integer Linear Program (MILP) formulation with unequal power-sharing for $N$ elements in (1) is used for comparison.
+Also, a Mixed-Integer Linear Program (MILP) formulation with both equal and unequal power-sharing for $N$ elements in (1) used for comparison.
 The battery element parameters given in Table I were used.
 The two use-cases are a) power reference tracking and b) revenue maximization.
+All cases are modeled using JuMP in Julia and were solved using Gurobi v12.0.0.
+All cases were run on a Macbook Pro with an M1 Pro processor.
 IV-A Power Reference Tracking
@@
 $\min_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K-1}\left((P_{\text{c}}[k]-P_{\text{d}}[k])-P_{\text{ref}}[k]\right)^{2}.$ (14)
-A timestep of $\Delta t=5$ min, along with $N=100$ was used.
+A timestep of $\Delta t=3$ min, along with $N=100$ was used.
 The RCB model was used three times with $M=1$, $M=5$, and $M=10$ respectively.
 The results from using each model are shown in Fig.
-5 and the solve times along with the mean squared tracking error are shown in Table II.
+4 and the solve times along with the mean squared error (MSE) are shown in Table II.
+Note that the relaxed model resulted in a dispatch with simultaneous charging and discharging, which cannot be realized under equal power sharing and results in the discrepancy between predicted MSE and actual MSE obtained while realizing the optimal dispatch.
+The net power trajectory was implemented which caused the BESS to saturate causing the discrepancy between the predicted and realized MSE values.
 Figure 4: Open-loop implementation of power optimal reference tracking power schedules using MILP, relaxed, robust and proposed RCB models.
-TABLE II: Power Reference Tracking Performance Metrics
+TABLE II: Power Reference Tracking Performance Metrics ($N=100$)
 Model Solver Predicted Actual
-Time (ms) MSE (kW2) MSE (kW2)
-MILP 50,140 0 0
-Relaxed 30 0 1841
-Robust 34 2.12 2.12
-RCB (M = 1) 30 0.72 0.72
-RCB (M = 5) 30 0.013 0.013
-RCB (M = 10) 30 0 0
-The MILP model yielded the optimal solution and was able to track the reference signal with no tracking error.
-However, the MILP also came with a significant computational burden compared to the rest of the formulations.
+Time MSE (kW2) MSE (kW2)
+MILP (unequal) 492.60 s 0 0
+MILP (equal) 34.37 s 200.0 200.0
+Relaxed 45 ms 0 1889
+Robust 54 ms 212 212
+RCB (M = 1) 50 ms 23.93 23.93
+RCB (M = 5) 50 ms 0.231 0.231
+RCB (M = 10) 50 ms 0 0
+The MILP model under unequal power sharing was able to track the reference signal with no tracking error.
+The MILP under equal power sharing couldn’t track the reference signal as well, despite both solving to 0% optimality gap, indicating that there is a benefit of simultaneously charging and discharging in this use case.
+The MILP, under both equal and unequal power sharing, came with a significant computational burden compared to the rest of the formulations, attributed to the extra binary variables added to enforce the complementarity constraint.
 The other formulations are linear and, as a result, have significantly faster solve times.
-The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs leading to a large realized tracking error.
+The relaxed model predicts zero tracking error; however, the solution was unrealizable and caused the battery to saturate around $t=6$ hrs, leading to a large realized tracking error.
 The robust model yields a realizable yet sub-optimal solution.
-The RCB model shows improved tracking error with increasing values of $M$ but even with $M=1$ the tracking error surpasses the robust model.
+The RCB model shows improved tracking error with increasing values of $M$, but even surpassing the robust model with $M=1$.
 IV-B Revenue Maximization
@@
 $\max_{P_{\text{c}}[k],P_{\text{d}}[k]}\qquad\sum_{k=0}^{K}C[k](P_{\text{d}}[k]-P_{\text{c}}[k]),$ (15)
-where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices for 24 hours, plotted in Fig 6.
+where $C[k]$ is the price at time $k$. The price signal was obtained from the California Independent System Operator (CAISO) day-ahead market prices, plotted in Fig 5.
 Figure 5: CAISO day-ahead price forecasts
@@
 An additional case was added with $M=900$ ($\delta t=1s$) to gain insight on model performance as $M\rightarrow\infty.$The results for the different models are shown in Table III.
-TABLE III: Revenue Maximization Performance metrics
-Model Solver Predicted Actual
-Time (ms) Revenue (\$) Revenue (\$)
-MILP 1250 2088.97 2088.97
-Relaxed 1 2092.80 1940.35
-Robust 1 1866.42 1866.42
-RCB (M = 1) 1 1411.53 1411.53
-RCB (M = 5) 1 1955.76 1955.76
-RCB (M = 10) 1 2022.87 2022.87
-RCB (M = 900) 1 2089.23 2089.23
-Similar to the previous use-case, the MILP model produced the highest revenue but at a significantly larger computational cost.
-The relaxed model predicts the highest revenue but underestimates the SOC, leading to unpredicted battery saturation and a realized performance that is significantly worse.
+TABLE III: Revenue Maximization Performance metrics ($N=100$)
+Model Solver Predicted Actual Optimality
+Time Revenue (\$) Revenue (\$) Gap (%)
+MILP (unequal) $90$ mins 2088.99 2088.99 0.02
+MILP (equal) 23 mins 2088.97 2088.97 0
+Relaxed 1.5 ms 2092.80 1940.35 0
+Robust 2.2 ms 1866.42 1866.42 0
+RCB (M = 1) 1.7 ms 1642.24 1642.24 0
+RCB (M = 5) 1.9 ms 2000.50 2000.50 0
+RCB (M = 10) 1.7 ms 2045.24 2045.24 0
+RCB (M = 900) 1.6 ms 2089.48 2089.48 0
+The MILP models produced high revenues that could also be realized, but at substantial computational cost, with unequal power sharing taking significantly longer to solve and was not able to solve to 0% optimality gap.
+The relaxed model predicts the highest revenue but underestimates the SOE, leading to unpredicted battery saturation and a realized performance that is significantly worse.
 The robust model yields a realizable yet suboptimal solution.
 The RCB model for small values of $M$ results in a sub-optimal solution which can be attributed to the larger timestep used in this use-case leading to a larger value or $\varepsilon$ and, in turn, a more restricted feasible set.
-However, for f $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and approaches the predicted performance for larger values of $M$.
-It was also observed that the RCB surpasses the robust model for $M\geq 3$ ($\delta t\leq 5$ min).
+However, for $M\geq 5$ ($\delta t\leq 3$ min), the RCB model surpasses all other linear models and for larger values of $M$ the predicted revenue steadily increases, maintaining realizability.
+We note that for $M=900$ the revenue even surpasses the element-wise MILP solution, because individual elements are able to switch from charging to discharging and back at a faster controller timestep when implemented by the PSC.
+A MILP solution with the control time step $\delta t$ would again achieve or exceed this result, but at much larger computational cost.
+The RCB model captures what is feasible at the fast control time step $\delta t$ without added complexity.
 V Conclusion
-A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements to realize simultaneous charging and discharging.
+A new linear dispatch model is presented and analyzed for a composite battery that permits unequal power sharing among its $N$ elements.
 A priority-based disaggregation strategy is described and used to characterize sufficient conditions on the composite battery dispatch under which element constraints are guaranteed to be satisfied.
@@
 By enabling unequal power sharing within a linear formulation, we have expanded the class of battery models that yield realizable power trajectories.
-Practical implementation of such dispatch strategies could also justify the use of linear models for composite storage systems or aggregate storage representations in energy system optimization models.
+Implementing the proposed controller is of interest and mainly requires communications to/from battery elements faster than $\delta t$ (similar to VPPs in the field today).
+Furthermore, embedding the presented linear models into energy system planning problems would inform the impact of battery systems more broadly.
+Future work involves exploring other disaggregation strategies and non-linear battery models.
 […]